OpenStudy (anonymous):

At the moment, Boris Holdings owns eleven more properties than A.C Realty. Last year Boris owned twice as many as they do now, and A.C three times as many. At that time Boris owned ten more than A.C. How many properties do Boris Holdings and A.C Realty currently own?

OpenStudy (anonymous):

let no. owned by AC be x, boris be y: x+11=y 3x+10 = 2y solving it simaltaneously: 3x+10=2y 2x+22=2y x-12=0, x=12 12+11=y=23 AC owns 12 Boris owns 23

OpenStudy (anonymous):

Thank you